if your job hires 200 people and 170 are paid hourly and the other 30 salary what percentage of the ppl are being paid hourly

Answer by stanbon(75887)

if your job hires 200 people and 170 are paid hourly and the other 30 salary what percentage of the ppl are being paid hourly
---------
x/100 = 170/200
---
x = 100(170/200)
----
x = 170/2
---
x = 85
------
Ans: 85%
===============
Cheers,
Stan H.